# Cross-Sectional Dynamics Under Network Structure: Theory and Macroeconomic Applications

### v6

@@ 2.1 Bilateral Connections in Networks
A network is represented by an $n\times n$ adjacency matrix $A$ with elements $a_{ij}$. I consider a directed, weighted and possibly signed network; $a_{ij}\in[-1,1]$ shows the sign and strength of the link from cross-sectional unit $i$ to unit $j$, with $a_{ij}\neq a_{ji}$ possibly. If $a_{ij}=0$, I say unit $i$ is not connected to unit $j$. Self-links $a_{ii}\neq 0$ are permitted. The set of bilateral links $\{a_{ij}\}_{i,j=1:n}$ give rise to a myriad of higher-order (hyper-dyadic) links among units, referred to as walks. [^13] Whenever convenient to simplify notation, I write $a:b$ for the set of integers $\{a,a+1,...,b\}$, $a\leq b$.
@@
A walk is the product of bilateral links that lead from unit $i$ to unit $j$ over some intermediary units. It is non-zero if all of these units are sequentially connected. Simple matrix algebra reveals that $A^{K}]_{ij}$ contains the sum of walks from $i$ to $j$ of length $K$. For $K\geq 1$, I refer to this quantity as the $K$th-order connection from $i$ to jj.141414 For $K=1$, the first-order connection $A^{1}]_{ij}=a_{ij}$ boils down to the direct link from $i$ to $j$. Consider the following example:
@@ 2.2 Lagged Innovation Transmission via Bilateral Links
Consider a stationary cross-sectional time series $y_{t}=(y_{1t},...,y_{nt})^{\prime}$ with mean zero. The NVAR formalizes the idea by which the dynamics of $y_{t}$ are driven by the lagged transmission of innovations $u_{it}$ along bilateral links $a_{ij}$ among cross-sectional units $i,j=1:n$. Throughout, I assume that links are fixed and that transmission is uni-directional: the direct link from $i$ to $j$, $a_{ij}$, transmits innovations from $j$ to $i$. Innovations may be cross-sectionally correlated.
@@ 2.2.1 Single Lag in Innovation Transmission: NVAR(1,1)
Consider the following VAR(1):
$y_{t}=\Phi y_{t-1}+u_{t}\;,\quad u_{t}\sim WN(0,\Sigma)\;,$ (1)
and let $\Phi=\alpha A$ for $\alpha\in\mathbb{R}$ be proportional to the adjacency matrix $A$. Under this process, the one step-ahead expectation of $y_{it}$ is proportional to the weighted sum of one period-lagged values of $y_{jt}$ for all units $j$ to which $i$ is directly linked, with weights given by links $\{a_{ij}\}_{j=1:n}$: $\mathbb{E}_{t-1}[y_{it}]=\alpha\sum_{j=1}^{n}a_{ij}y_{j,t-1}$. For example, in the RBC input-output economy of Long and Plosser (1983), the expected output in sector $i$ tomorrow is a weighted average of the production in its supplier-sectors $j$ today. In Golub and Jackson (2010), individual $i$’s expected opinion tomorrow is a weighted average of their friends’ opinions today. For reasons clarified below, I dub this process NVAR(1,1).
Under Eq. 1, dynamics of $y_{t}$ at horizon $h\geq 1$ are driven by $h$th-order network connections. More precisely, assuming $\alpha\neq 0$, $y_{j}$ Granger-causes $y_{i}$ at horizon $h$ iff there exists an $h$th-order connection from $i$ to jj:151515 In other words, given all other variables $y_{kt}$ for $k\neq j$, $y_{jt}$ is useful in forecasting $y_{i,t+h}$ iff there exists an $h$th-order connection from $i$ to $j$.
@@
Fig. 1 illustrates these Granger-causality dynamics – also referred to as Generalized Impulse-Response Functions (GIRFs)161616 The GIRF is “generalized” because it disregards shock identification, but considers the propagation of reduced form errors $u_{t}$ over future time periods $t+h$, $h>0$. – for the network from Section 2.1 under $\alpha=1$. It points to two insights. First, lagged network interactions generate persistence; even white noise-innovations $\{u_{jt}\}_{j=1}^{n}$ cause $y_{t}$ to react persistently because each $u_{jt}$ affects connected units $i\neq j$ only with a lag. [^17] Furthermore, each individual $y_{jt}$ reacts persistently to its own white noise-innovation $u_{jt}$ as long as there is a cycle in the network, i.e. a walk from $j$ to $j$. Notably, this can hold even without a self-link – $a_{jj}=0$. This result is behind the “endogenous business cycles” in Long and Plosser (1983)’s economy with lagged input-output conversion. Second, under this type of persistence, network-connections affect not only the strength of impulse-responses, but also their timing. For example, while unit 2 is directly linked to unit 1 and therefore experiences the latter’s innovation with a lag of one period, unit 3 only has a second-order connection to unit 1 and therefore reacts only after two periods. This result relates to Dufour and Renault (1998), who point out that Granger-causality can take the form of chains.181818 Specifically, even if a series $x_{t}$ does not Granger-cause a series $y_{t}$ at horizon 1, under the presence of a third series $z_{t}$, $x_{t}$ might Granger-cause $y_{t}$ at higher horizons as the causality could run from $x_{t}$ to $z_{t}$ to $y_{t}$. In the case of a (cross-sectional) time series driven by lagged innovation transmission along bilateral links, their insight emerges naturally and their generally non-trivial conditions for Granger-(non)causality boil down to the presence or absence of network connections of relevant order between the relevant variables (units). As I show in the following, this holds even under more general transmission-timing than considered in Eq. 1.
@@
While the process in Eq. 1 reveals useful theoretical insights in macro- (Long and Plosser, 1983) and microeconomics (Golub and Jackson, 2010) alike, it appears too restrictive for a range of empirical work. It assumes that innovations transmit at the speed of one link per period and that transmission materializes completely within a single period. This implies, for instance, that in response to news gathered by a friend of a friend, $j$, an individual $i$ adjusts their opinion in two periods, not earlier, not later. Similarly, an innovation to a sector $j$’s output affects the output of other sectors $i$ located two positions downstream (“customers of customers”) in exactly two periods. In either case, earlier responses are ruled out altogether, and later adjustments only occur if $i$ has a third- or higher-order connection to $j$. In the following, I generalize the timing of lagged network effects by extending the simple process above along two dimensions.
@@ 2.2.2 Multiple Lags in Innovation Transmission: NVAR($p,1$)
$y_{t}=\alpha_{1}Ay_{t-1}+...+\alpha_{p}Ay_{t-p}+u_{t}\;,\quad u_{t}\sim WN(0,\Sigma)\;,$ (2)
@@
Let $y_{t}$ evolve as in Eq. 2. Then, Granger-causality from $y_{j}$ to $y_{i}$ at horizon $h$ is a linear combination of connections from $i$ to $j$ of order $k\in\{\underline{k},\underline{k}+1,...,h\}$, where $\underline{k}=ceil(h/p)$.191919ceil(x)ceil(x) rounds $x\in\mathbb{Q}$ up to the next integer.
@@
Each coefficient $c^{h}_{k}$, for $k=\underline{k}:h$, is a polynomial of $\{\alpha_{l}\}_{l=1:p}$ and shows the importance of connection-order $k$ for the impulse-response at horizon $h$. If $i$ only has a first- but no higher-order connections to $j$, then Eq. 3 boils down to $\partial y_{i,t+h}/\partial u_{j,t}=\alpha_{h}a_{ij}$ for $h=1:p$ and equals zero otherwise. Hence, the NVAR($p,1$) specifies that dynamics are driven by the lagged transmission of innovations along direct links over $p$ periods. $\alpha$ determines how this transmission is spread out over these $p$ periods and, consequently, how transmission along higher-order connections accrues as time progresses.202020 Note that the elements of $\alpha$ are allowed to be negative. For example, under $p=2$, $(\alpha_{1}>0,\alpha_{2}<0)$ with $\alpha_{1}+\alpha_{2}>0$ signifies an initial overreaction and subsequent correction of $y_{it}$ after an innovation to some $y_{jt}$ s.t. $a_{ij}>0$. Note that the transmission is assumed to be the same for all unit pairs $(i,j)$ and invariant over time.
@@ 2.2.3 Multiple Rounds of Innovation Transmission: NVAR($p,q$)
I dub this process NVAR($p,q$). $q$ indicates the relation between the frequency of network interactions, at which $\tilde{y}_{\tau}$ evolves, and the frequency at which data $\{y_{t}\}_{t=1:T}$ is observed. If $q=1$, the two coincide, and, trivially, $y_{t}$ and $\tilde{y}_{\tau}$ are the same NVAR($p,1$) process. Instead, if $q>1$, then network interactions occur at higher frequency than data is observed; for example, under yearly observations, $q=4$ implies quarterly network interactions. As a result, multiple rounds of innovation transmission occur in a single observational period and dynamics at horizon $h$ are affected by higher-order connections:
@@
Let $y_{t}$ evolve as in Eq. 4 for some $q\in\mathbb{N}$. Then, Granger-causality from $\tilde{y}_{j}$ to $\tilde{y}_{i}$ at horizon $h$ is a linear combination of connections from $i$ to $j$ of order $k\in\{\underline{k},\underline{k}+1,...,hq\}$, where $\underline{k}=ceil(hq/p)$.212121 Assuming again $\alpha_{l}>0\;\forall\;l$ and $a_{ij}\geq 0$, Proposition 2 implies that $\tilde{y}_{j}$ Granger-causes $\tilde{y}_{i}$ at horizon $h$ iff there exists a connection from $i$ to $j$ of at least one order $k\in\{\underline{k},\underline{k}+1,...,hq\}$.
@@
with similar properties as under Eq. 4. See Section A.4.222222 See Proposition 10 and what follows in Section 2.3. Also, analogous calculations apply if $y_{t}=\tilde{y}_{\tau}+...+\tilde{y}_{\tau-q+1}$.
By restricting the NVAR($p,q$) in Eq. 4, we can accommodate more intricate relations between the frequencies of network interactions and observation; for any $q\in\mathbb{Q}_{++}$, we can write $q=q_{1}q_{2}$ with $q_{2},q_{1}^{-1}\in\mathbb{N}$, and we can model $y_{t}$ as an NVAR($p^{*},q_{2})$ with $p^{*}=p/q_{1}\in\mathbb{N}$ and with restricted parameters $\{\alpha_{l}\}_{l=1:p^{*}}$ to comply with the stated interpretations of $p$ and $q$. For example, under monthly observations, $q=4/3$ signifies that network interactions occur (roughly) every three weeks. This amounts to observing every $q_{2}=4$ periods a snapshot of a weekly process that depends on its value in the past $p$ three-week-periods, i.e. on its value three weeks ago, six weeks ago, and so forth, until $p^{3}=3p$ weeks ago. See Section A.4 for more details.232323 If $y_{t}$ is observed for $T$ periods – $t=1:T$ –, then $\tilde{y}_{\tau}$ evolves for $T_{\tau}$ (network interaction-)periods – $\tau=1:T_{\tau}$. Thereby, $T_{\tau}$ satisfies $T=|\{(1:T_{\tau})/q\}\cap\mathbb{N}|$; the number of elements in the set $1:T_{\tau}$ that are integer-multiples of $q$ shall equal $T$. This yields $T_{\tau}=qT$ under $q\in\mathbb{N}$, and $T_{\tau}=T$ under $q^{-1}\in\mathbb{N}$. For other $q\in\mathbb{Q}_{++}$, we have $T_{\tau}=q_{2}T$, where $q_{2}$ is least common multiple of $q$ and $1$.
@@ 2.3 Further Properties of the NVAR
Corollary 1 characterizes stationarity of the NVAR($p,q$) in terms of eigenvalues of the network adjacency matrix $A$ and the time profile of network effects $\alpha$.242424 Corollary 1 follows from Propositions 8, 9 and 13. Intuitively, the conditions ensure that $\lim_{k\rightarrow\infty}\;c(\alpha,k)A^{k}=0$, for any polynomial in $\alpha$ of order $k$, $c(\alpha,k)$, which by Proposition 1 ensures vanishing effects of disturbances. Under $p>1$, and ever more so the larger $n$, this drastically simplifies checking for stationarity. Note that the second statement amounts to weak stationarity of the AR($p$) process $(1-\lambda_{i}\alpha_{1}L-...-\lambda_{i}\alpha_{p}L^{p})x_{t}=v_{t}$ for each eigenvalue $\lambda_{i}$ of $A$, though AR-coefficients are typically required to be real.
@@
Provided that $|\lambda_{i}|<1$ for all eigenvalues $\lambda_{i}$ of $A$, we can write $x=(I-A)^{-1}v=(A+A^{2}+A^{3}+...)v$. Hence, the implicit assumption is that connections of all order materialize in a single period. In line with that, Corollary 2 establishes that the response of $x_{i}$ to an innovation $v_{j}$ in the SAR model from Eq. 6 is equal to the long-run response of $y_{it}$ to a persistent innovation $\tilde{u}_{j\tau}$ in the NVAR($p,q$) from Eqs. 4 and 5.262626 Corollary 2 follows from Proposition 5 and Proposition 11.
@@
Both responses are given by the Leontief inverse $(I-aA)^{-1}$, which is a sufficient statistic for the long-term cross-sectional comovement. By taking a stance on the timing and frequency of network interactions, the NVAR shows how any such long-term effect materializes over time, which allows the researcher to go beyond steady state comparisons and characterize transitional dynamics.272727 Note the correspondence between the timing of transitional dynamics and the timing of responses to temporary innovations; as for any VAR, the long-term response to a permanent innovation is equal to the cumulative response to temporary innovations.
The long-run is defined in terms of the frequency of network interactions. To justify the use of the SAR at the expense of the NVAR, then, we need this frequency to be high- and the data frequency to be low enough. Corollary 3 formalizes this idea; if we define the observed process $y_{t}$ as the sum of an underlying process evolving at the higher frequency of network interactions, then, as this frequency diverges to infinity, $y_{t}$’s response to a high-frequency innovation that occurs within observational period $t$ converges to the response of $x_{i}$ to an innovation $v_{j}$ in the SAR model from Eq. 6.282828 It follows from Proposition 5 and Proposition 12, applied for $\rho_{j}=0$. Thus, an SAR for $x$ can be derived by writing $x$ as the sum of an underlying process driven by lagged network interactions that evolves at an ever higher frequency.
@@
$\underset{q,G\rightarrow\infty}{\lim}q\frac{\partial y_{t}}{\partial\tilde{u}_{tq-G}}=(I-aA)^{-1}\;.$
@@
Instead of relying on lagged network interactions, one can also introduce transitional dynamics to the SAR model from Eq. 6 through autocorrelated innovations $v$. This approach is predominant in macroeconomic studies of input-output-economies; under fairly standard assumptions (see e.g. Carvalho and Tahbaz-Salehi (2019)), sectoral production and prices evolve according to Eq. 6, whereby $v$ is Total Factor Productivity (TFP), which – across all areas of macroeconomics – is typically assumed to follow an autoregressive process.292929 It can be split into an idiosyncratic and an aggregate component; see Section 4.
Lagged network interactions and autocorrelated innovations lead to different kinds of dynamics. As first illustrated with the NVAR(1,1) and Fig. 1, and as discussed throughout this section for the more general NVAR($p,q$), lagged network interactions relate impulse-responses to network-connections of different order, whereby the precise mapping is determined by $\alpha$ and $q$. As a consequence, how strongly $y_{i}$ reacts to an impulse to $y_{j}$ depends on the strength of network-connections from $i$ to $j$ of relevant order, and how fast it reacts depends on how many lower- as opposed to higher-order connections the two units share. For instance, in the environment of Golub and Jackson (2010) this implies that an individual adjusts their opinion faster after news shared by closer than more distant friends. Similarly, in the context of Long and Plosser (1983) it means that a sector contracts its production faster after a negative productivity shock to more immediate suppliers than to ones located further upstream.
@@
The strength of connections from $i$ to $j$ merely scales the impulse-response; network connections merely amplify exogenous dynamics.
These insights carry over to the case when the two sources of dynamics are combined. Consider an NVAR($p,1$) for simplicity, and let $(1-\rho_{j}L)u_{jt}=\varepsilon_{jt}$. Then, using Proposition 1,
@@
Fig. 2 illustrates for an NVAR($2,1$) and $\rho_{j}=.6$, once under $\alpha=(.8,.2)^{\prime}$ (top row), once under $\alpha=(.2,.8)^{\prime}$ (bottom row). The left panel depicts how connection-orders accumulate over horizons under such $\alpha$, while the middle and right panel show the resulting impulse-responses for two sectors that share only a first- or only a second-order connection, respectively. Under $\alpha=(.8,.2)^{\prime}$, the innovation travels faster through the network than under $\alpha=(.2,.8)^{\prime}$. In the former case, the transmission along a direct link (darkest blue dots and bars) is strongest after one period, in the latter after two periods. Consequently, in the former case, the transmission along second-order connections (lighter blue dots and bars) is strongest in the second period, in the latter in the fourth period. The exogenous persistence ensures that the fraction $\rho_{j}$ of any transmission at horizon $h$ is carried over to horizon $h+1$, a fraction $\rho_{j}^{2}$ to horizon $h+2$, etc. (ever lighter gray bars).
@@
Even under autocorrelated innovations, lagged and contemporaneous network interactions yield the same long-run responses. This is shown by Corollary 4.303030 Corollary 4 follows from Proposition 6 and Proposition 11. In contrast to Corollary 2, we consider the long-run rather than contemporaneous response of $x$, since the implicit assumption in the SAR model with autocorrelated innovations is that they evolve at observational frequency (while the frequency of network interactions diverges to infinity).

### v7

@@ 2.1 Bilateral Connections in Networks
A network is represented by an $n\times n$ adjacency matrix $A$ with elements $a_{ij}$. I consider a directed, weighted and possibly signed network; $a_{ij}\in[-1,1]$ shows the sign and strength of the link from unit $i$ to unit $j$, with $a_{ij}\neq a_{ji}$ possibly. If $a_{ij}=0$, I say unit $i$ is not connected to unit $j$. Self-links are permitted: $a_{ii}\neq 0$. The set of bilateral links $\{a_{ij}\}_{i,j=1:n}$ give rise to a myriad of higher-order (hyper-dyadic) links among units, referred to as walks. [^7] Whenever convenient to simplify notation, I write $a:b$ for the set of integers $\{a,a+1,...,b\}$, $a\leq b$.
@@
A walk is a product of bilateral links that lead from unit $i$ to unit $j$ over intermediary units. It is non-zero if all of these units are sequentially connected. Simple matrix algebra reveals that $A^{K}]_{ij}$ contains the sum of walks from $i$ to $j$ of length $K$. I refer to this quantity as the $K$th-order connection from $i$ to $j$. Consider the following example:
@@ 2.2 Lagged Innovation Transmission via Bilateral Links
Consider a stationary cross-sectional time series $y_{t}=(y_{1t},...,y_{nt})^{\prime}$ with mean zero. Under an NVAR, the dynamics of $y_{t}$ are driven by the lagged transmission of innovations $u_{it}$ along bilateral links $a_{ij}$ among cross-sectional units $i,j=1:n$. Throughout, I assume that links are fixed over time and that transmission is uni-directional: the direct link from $i$ to $j$, $a_{ij}$, transmits innovations from $j$ to $i$. Innovations $u_{t}$ may be cross-sectionally correlated.
@@ 2.2.1 Single Lag in Innovation Transmission: NVAR(1,1)
Consider a VAR(1) with an autoregressive matrix proportional to the adjacency matrix $A$:
$y_{t}=\alpha Ay_{t-1}+u_{t}\;,$ (1)
for $\alpha\in\mathbb{R}$. Under this process, the one step-ahead expectation of $y_{it}$ is proportional to the weighted sum of one period-lagged values of $y_{jt}$ for all units $j$ to which $i$ is directly linked: $\mathbb{E}_{t-1}[y_{it}]=\alpha\sum_{j=1}^{n}a_{ij}y_{j,t-1}$. For example, in the RBC input-output economy of Long and Plosser (1983), the expected output in sector $i$ tomorrow is a weighted average of the production in its supplier-sectors $j$ today. In Golub and Jackson (2010), individual $i$’s expected opinion tomorrow is a weighted average of their friends’ opinions today. For reasons clarified below, I dub this process NVAR(1,1).
Under Eq. 1, dynamics of $y_{t}$ at horizon $h\geq 1$ are driven by $h$th-order network connections. More precisely, assuming $\alpha\neq 0$, $y_{j}$ Granger-causes $y_{i}$ at horizon $h$ iff there exists an $h$th-order connection from $i$ to jj: [^8] In other words, given all other variables $y_{kt}$ for $k\neq j$, $y_{jt}$ is useful in forecasting $y_{i,t+h}$ iff there exists an $h$th-order connection from $i$ to $j$.
@@
Fig. 1 illustrates these Granger-causality dynamics – also referred to as Generalized Impulse-Response Functions (GIRFs) [^9] The GIRF is “generalized” because it disregards shock identification, but considers the propagation of reduced form errors $u_{t}$ over future time periods $t+h$, $h\geq 1$. – for the network from Section 2.1 under $\alpha=1$. It points to two insights. First, lagged network interactions generate persistence; even white noise-innovations $\{u_{jt}\}_{j=1}^{n}$ lead to persistent reactions of $y_{t}$ because each $u_{jt}$ affects connected units $i\neq j$ with a lag. [^10] Furthermore, each individual $y_{jt}$ reacts persistently to its own white noise-innovation $u_{jt}$ as long as there is a cycle in the network, i.e. a walk from $j$ to $j$. Notably, this can hold even without a self-link – $a_{jj}=0$. This result is behind the endogenous business cycles in Long and Plosser (1983)’s economy with lagged input-output conversion. Second, under this type of persistence, network-connections affect not only the strength of impulse-responses, but also their timing. For example, while unit 2 is directly linked to unit 1 and therefore experiences the latter’s innovation with a lag of one period, unit 3 only has a second-order connection to unit 1 and therefore reacts only after two periods. This result relates to Dufour and Renault (1998), who point out that Granger-causality can take the form of chains. [^11] Specifically, even if a series $x_{t}$ does not Granger-cause a series $y_{t}$ at horizon 1, under the presence of a third series $z_{t}$, $x_{t}$ might Granger-cause $y_{t}$ at higher horizons as the causality could run from $x_{t}$ to $z_{t}$ to $y_{t}$. In the case of a (cross-sectional) time series driven by lagged innovation transmission along bilateral links, their insight emerges naturally and their generally non-trivial conditions for Granger-(non)causality boil down to the presence or absence of network connections of relevant order between the relevant variables (units). As I show in the following, this holds even under a more general transmission-timing than considered in Eq. 1.
@@
While the process in Eq. 1 reveals useful theoretical insights in both macro- (Long and Plosser, 1983) and microeconomics (Golub and Jackson, 2010), it is too restrictive for most empirical work. It assumes that innovations transmit at the speed of one link per period and that transmission materializes completely within a single period. This implies, for instance, that in response to news gathered by a friend of a friend, $j$, an individual $i$ adjusts their opinion in two periods, not earlier, not later. Similarly, an innovation to a sector $j$’s output affects the output of other sectors $i$ located two positions downstream (“customers of customers”) in exactly two periods. In either case, earlier responses are ruled out altogether, and later adjustments only occur if $i$ has a third- or higher-order connection to $j$. In the following, I generalize the timing of lagged network effects by extending the simple process above along two dimensions.
@@ 2.2.2 Multiple Lags in Innovation Transmission: NVAR($p,1$)
$y_{t}=\alpha_{1}Ay_{t-1}+...+\alpha_{p}Ay_{t-p}+u_{t}\;,$ (2)
@@
Let $y_{t}$ evolve as in Eq. 2. Then, Granger-causality from $y_{j}$ to $y_{i}$ at horizon $h$ is a linear combination of connections from $i$ to $j$ of order $k\in\{\underline{k},\underline{k}+1,...,h\}$, where $\underline{k}=ceil(h/p)$.121212ceil(x)ceil(x) rounds $x\in\mathbb{Q}$ up to the next integer.
@@
Each coefficient $c^{h}_{k}$ for $k=\underline{k}:h$ is a polynomial of $\{\alpha_{l}\}_{l=1:p}$ and shows the importance of connection-order $k$ for the impulse-response at horizon $h$. If $i$ has a first- and no higher-order connections to $j$, then $\partial y_{i,t+h}/\partial u_{j,t}=\alpha_{h}a_{ij}$ for $h=1:p$ and zero otherwise. Hence, the NVAR($p,1$) specifies that dynamics are driven by the transmission of innovations along direct links lagged over $p$ periods. $\alpha$ determines how the transmission is spread out over these $p$ periods and, consequently, how transmission along higher-order connections accrues as time progresses. Note that the transmission is assumed to be the same for all unit pairs $(i,j)$ and invariant over time.
88todo: 8note that $\alpha\neq$ timing of network effects, as $q$ matters too!
@@ 2.2.3 Multiple Rounds of Innovation Transmission: NVAR($p,q$)
I dub this process NVAR($p,q$). $q$ indicates the relation between the frequency of network interactions, at which $\tilde{y}_{\tau}$ evolves, and the frequency at which data $\{y_{t}\}_{t=1:T}$ is observed. If $q=1$, the two coincide and, trivially, $y_{t}$ and $\tilde{y}_{\tau}$ are the same NVAR($p,1$) process. Instead, if $q>1$, then network interactions occur at higher frequency than data is observed. For example, under yearly observations, $q=4$ implies quarterly network interactions. As a result, multiple rounds of transmission occur in a single observational period and dynamics at horizon $h$ are affected by higher-order connections, $k>h$:
@@
Let $y_{t}$ evolve as in Eq. 4 for some $q\in\mathbb{N}$. Then, Granger-causality from $\tilde{y}_{j}$ to $\tilde{y}_{i}$ at horizon $h$ is a linear combination of connections from $i$ to $j$ of order $k\in\{\underline{k},\underline{k}+1,...,hq\}$, where $\underline{k}=ceil(hq/p)$. [^13] Assuming again $\alpha_{l}>0\;\forall\;l$ and $a_{ij}\geq 0$, Proposition 2 implies that $\tilde{y}_{j}$ Granger-causes $\tilde{y}_{i}$ at horizon $h$ iff there exists a connection from $i$ to $j$ of at least one order $k\in\{\underline{k},\underline{k}+1,...,hq\}$.
@@
with similar properties as under Eq. 4.141414See Section A.4 and Section 2.3. Analogous calculations apply if $y_{t}=\tilde{y}_{\tau}+...+\tilde{y}_{\tau-q+1}$.
By restricting the NVAR($p,q$) in Eq. 4, we can accommodate more intricate relations between the frequencies of network interactions and observation; for any $q\in\mathbb{Q}_{++}$, we can write $q=q_{1}q_{2}$ with $q_{2},q_{1}^{-1}\in\mathbb{N}$, and we can model $y_{t}$ as an NVAR($p^{*},q_{2})$ with $p^{*}=p/q_{1}\in\mathbb{N}$ and with restricted parameters $\{\alpha_{l}\}_{l=1:p^{*}}$ to comply with the stated interpretations of $p$ and $q$. For example, under monthly observations, $q=4/3$ signifies that network interactions occur (roughly) every three weeks. This amounts to observing every $4$ periods a snapshot of a weekly process that depends on its value in the past pp 33-week-periods, i.e. on its value three weeks ago, six weeks ago, and so forth, until $p^{*}=3p$ weeks ago. Section A.4 elaborates.
@@ 2.3 Further Properties of the NVAR
In the NVAR, the interdependence of $y_{it}$ and $y_{j,t-h}$ is an interplay of the cross-sectional distance between units $i$ and $j$ – encoded by the network – and the temporal distance between periods $t$ and tht-h. Correspondingly, Corollary 1 characterizes stationarity in terms of eigenvalues of the network adjacency matrix $A$ and roots of an AR process shaped by the timing of innovation transmission along a single link, $\alpha$. [^15] Corollary 1 follows from Propositions 8, 9 and 13. Intuitively, the conditions ensure that $\lim_{k\rightarrow\infty}\;c(\alpha,k)A^{k}=0$, for any polynomial in $\alpha$ of order $k$, $c(\alpha,k)$. By Proposition 1, this ensures that the effects of disturbances vanish for higher horizons. Under $p>1$, this simplifies checking for stationarity, especially for large $n$. Note that the second statement requires the AR($p$) process $(1-\phi_{1}L-...-\phi_{p}L^{p})x_{t}=v_{t}$ with $\phi_{l}=\alpha_{l}\lambda_{i}$ to be stationary for each eigenvalue $\lambda_{i}$ of $A$, though AR-coefficients are typically required to be real-valued.
@@
Provided that $|\lambda_{i}|<1$ for all eigenvalues $\lambda_{i}$ of $A$, we can write $x=(I-A)^{-1}v=(A+A^{2}+A^{3}+...)v$. Hence, the implicit assumption is that connections of all order materialize in a single period. In line with that, Corollary 2 establishes that the response of $x_{i}$ to an innovation $v_{j}$ in the SAR model from Eq. 6 is equal to the long-run response of $y_{it}$ to a persistent innovation $\tilde{u}_{j\tau}$ in the NVAR($p,q$) from Eq. 4 or Eq. 5.161616Corollary 2 follows from Propositions 5 and 11.
@@
Both responses are given by the Leontief inverse $(I-aA)^{-1}$, which is a sufficient statistic for the long-term cross-sectional comovement. By taking a stance on the timing and frequency of network interactions, the NVAR shows how any such long-term effects materialize over time; it goes beyond steady state comparisons and characterizes transitional dynamics. [^17] Note the correspondence between the timing of transitional dynamics and the timing of responses to temporary innovations; as for any VAR, the long-term response to a permanent innovation is equal to the cumulative response to temporary innovations. $\alpha$ temporally decomposes the spatial autoregressive parameter $a$.
The long-run is defined in terms of the frequency of network interactions. To justify the use of the SAR at the expense of the NVAR, then, we need this frequency to be high enough and the data frequency to be low enough. Corollary 3 formalizes this idea; if we define the observed process $y_{t}$ as the sum of an underlying process $\tilde{y}_{\tau}$ evolving at the higher frequency of network interactions, then, as this frequency diverges to infinity, $y_{t}$’s response to a high-frequency innovation that occurs within observational period $t$ converges to the response of $x_{i}$ to an innovation $v_{j}$ in the SAR model from Eq. 6.181818Corollary 3 follows from Propositions 5 and 12, applied for $\rho_{j}=0$.
@@
$\underset{g\rightarrow\infty}{\lim}q\frac{\partial y_{t}}{\partial\tilde{u}_{\tau-g}}=(I-aA)^{-1}\;.$
@@
Instead of relying on lagged network interactions, one can also introduce transitional dynamics to the SAR model from Eq. 6 through autocorrelated innovations $v$. This approach is predominant in macroeconomic studies of input-output-economies; under fairly standard assumptions (see e.g. Carvalho and Tahbaz-Salehi (2019)), sectoral production and prices evolve according to Eq. 6, whereby $v$ is Total Factor Productivity (TFP), which – across all areas of macroeconomics – is typically assumed to follow an autoregressive process. [^19] It can be split into an idiosyncratic and an aggregate component; see Section 4. 99todo: 9and in SAR some allow for autocorr. errors?1010todo: 10also, cite some others in footnote on idiosync. vs. agg. component
Lagged network interactions and autocorrelated innovations lead to different kinds of dynamics. As first illustrated in Fig. 1 for the NVAR(1,1), and as discussed throughout this section for the general NVAR($p,q$), lagged network interactions relate impulse-responses to network-connections of different order, whereby the precise mapping is determined by $\alpha$ and $q$. As a consequence, how strongly $y_{i}$ reacts to an impulse to $y_{j}$ depends on the strength of network-connections from $i$ to $j$ of relevant order, and how fast it reacts depends on how many lower- as opposed to higher-order connections the two units share. For instance, in the environment of Golub and Jackson (2010) this implies that an individual adjusts their opinion faster after news shared by closer than more distant friends. Similarly, in the context of Long and Plosser (1983) it means that a sector contracts its production faster after a negative productivity shock to more immediate suppliers than to ones located further upstream.
@@
The strength of connections from $i$ to $j$ merely scales the exponentially decaying impulse-response; network connections merely amplify exogenous dynamics.
When the two sources of dynamics are combined, their qualitatively distinct contributions remain. Consider an NVAR($p,1$), and let $(1-\rho_{j}L)u_{jt}=\varepsilon_{jt}$. Then, using Proposition 1,
@@
Fig. 2 illustrates for an NVAR($2,1$) and $\rho_{j}=.6$, once under $\alpha=(.8,.2)^{\prime}$ (top row), once under $\alpha=(.2,.8)^{\prime}$ (bottom row). The left panel depicts how connection-orders accumulate over horizons for these $\alpha$, while the middle and right panels show the resulting impulse-responses for two sectors that share only a first- or only a second-order connection, respectively. Under $\alpha=(.8,.2)^{\prime}$, the innovation travels faster through the network than under $\alpha=(.2,.8)^{\prime}$; in the former case, the transmission along a direct link (darkest blue dots and bars) is strongest after one period, in the latter after two periods. Consequently, under $\alpha=(.8,.2)^{\prime}$, the transmission along second-order connections (lighter blue dots and bars) is strongest in the second period, under $\alpha=(.2,.8)^{\prime}$ in the fourth period. The exogenous persistence ensures that the fraction $\rho_{j}$ of any transmission at horizon $h$ is carried over to horizon $h+1$, a fraction $\rho_{j}^{2}$ to horizon $h+2$, etc. (ever lighter gray bars).
@@
Even under autocorrelated innovations, lagged and contemporaneous network interactions yield the same long-run responses. This is shown by Corollary 4.202020Corollary 4 follows from Propositions 6 and 11. In contrast to Corollary 2, it considers the long-run rather than contemporaneous response of $x$, since the implicit assumption in the SAR model with autocorrelated innovations is that they evolve at observational frequency.